# The Sound of an Orbit: A Quantum Spectrum at the ISCO

### v1

@@ I Introduction
The synthesis of quantum mechanics and general relativity remains one of the most profound challenges in modern theoretical physics Rovelli (2004). In this pursuit, the study of quantum fields in curved spacetimes has emerged as an indispensable tool, offering a window into the fundamental interactions that govern our universe Birrell and Davies (1982). Black holes, as cosmic laboratories of extreme gravity, provide a unique arena to probe the intricate link between quantum phenomena and the fabric of spacetime itself Bardeen et al. (1973). The behavior of quantum fields in the vicinity of these enigmatic objects not only deepens our understanding of foundational physics but also holds direct relevance for astrophysical observables, such as the celebrated Hawking radiation and the complex dynamics of accretion disks Page and Thorne (1974); Thorne (1974); Frank et al. (2002).
@@
A particularly fascinating frontier in this domain is the interaction between localized quantum systems, such as atoms, and the quantum fields permeating the curved background of a black hole. When an atom is subjected to a strong gravitational field, its internal energy levels are perturbed not only by the spacetime curvature but also by the incessant fluctuations of the quantum vacuum Audretsch and Muller (1995). This interaction can lead to observable phenomena, providing a novel way to probe the structure of both the gravitational field and the quantum vacuum itself. The theoretical underpinnings for such investigations are well-established, rooted in the Unruh effect, which predicts that an accelerated observer in flat spacetime will perceive the vacuum as a thermal bath, and its gravitational analogue, Hawking radiation, where black holes emit thermal particles due to quantum effects near their event horizon Hawking (1975); Unruh (1976).
@@
Building on these foundational concepts, previous studies have extensively examined the response of Unruh-DeWitt detectors, which are simplified yet powerful models for atoms, on various trajectories, including those held at static positions or undergoing radial infall toward a black hole. More recently, a quantum optics approach has been successfully applied to investigate the ”horizon brightened acceleration radiation” (HBAR) emitted by atoms falling into a black hole, revealing a deep connection between acceleration radiation, thermodynamics, and near-horizon conformal symmetries Scully et al. (2018); Azizi et al. (2021). This framework has proven to be a versatile tool for analyzing such complex quantum interactions.
@@
While these studies have provided invaluable insights, the specific case of a quantum detector following a stable circular geodesic has received comparatively less attention Garay et al. (2000); Louko and Satz (2006); Casals et al. (2009); Hodgkinson et al. (2014); Ng et al. (2014); Biermann et al. (2020). This paper seeks to fill this gap by focusing on a region of profound astrophysical and theoretical importance: the ISCO. Located at a radius of ($r=6M$) for a Schwarzschild black hole, the ISCO represents the final frontier of stable orbital motion, a critical boundary where matter in accretion disks accumulates before its final plunge. Understanding quantum interactions at this precipice is multifaceted Ottewill and Winstanley (2000). It promises to shed light on how quantum systems behave under extreme gravitational acceleration, potentially revealing novel quantum-gravitational effects. Furthermore, an atom orbiting at the ISCO can act as a sensitive probe of the quantum vacuum structure, offering a way to study effects akin to Hawking radiation without the need for direct observation of the horizon itself Barbado et al. (2011).
@@ II.1 Schwarzschild Spacetime and Geodesics at the ISCO
The gravitational field of a non-rotating, uncharged, spherically symmetric black hole of mass $M$ is described by the Schwarzschild metric. In geometric units, where the gravitational constant $G$ and the speed of light $c$ are set to unity, the line element in Schwarzschild coordinates $(t,r,\theta,\varphi)$ is given by Carroll (2019)
@@ II.2 The Unruh-DeWitt Detector and the Quantum Field
where $\sigma^{+}=|e\rangle\langle g|$ and $\sigma^{-}=|g\rangle\langle e|$ are the standard atomic raising and lowering operators, respectively. The interaction between the detector and a massless scalar field $\phi$ is described by the interaction Hamiltonian, written in the interaction picture as Unruh (1976); DeWitt (1979)
@@
[…]
@@
The choice of vacuum state, $|0\rangle$, is critical as it defines the initial state of the field with which the atom interacts. For an observer located at asymptotic infinity, the natural vacuum state is the Boulware vacuum, denoted $|0_{\text{B}}\rangle$. This state is defined by the condition that it is annihilated by the operators $a_{\omega lm}$ corresponding to mode functions $u_{\omega lm}$ that are purely positive frequency with respect to the global timelike Killing vector $\partial_{t}$. Physically, the Boulware vacuum corresponds to a state with no particles and no radiation as seen by a static observer far from the black hole. This is the ideal choice for our analysis, as it allows us to isolate the excitation of the atom that is due solely to its orbital motion, ensuring that our results are not conflated with the thermal bath of particles associated with the Hawking effect, which is characteristic of other vacuum states like the Hartle-Hawking vacuum.
@@ III.1 The Wightman Function along the ISCO Trajectory
The transition rate $R(\Omega_{0})$ for a detector with energy gap $\Omega_{0}$ is given by Fermi’s Golden Rule Brown (1992). For a stationary trajectory, such as a circular orbit, it is expressed as the Fourier transform of the Wightman function with respect to the proper time difference $\Delta\tau=\tau-\tau^{\prime}$:
@@
[…]
@@
Here, $\psi_{\omega l}(r)$ are the radial functions that solve the Regge-Wheeler equation, and $Y_{lm}(\theta,\varphi)$ are the standard spherical harmonics.
Our task is to evaluate the Wightman function specifically along the atom’s ISCO worldline, $x^{\mu}(\tau)=(\sqrt{2}\tau,6M,\pi/2,\tau/6\sqrt{3}M)$. Substituting these coordinates into the mode functions, we obtain the Wightman function as experienced by the orbiting atom as
[…]
×$\times e^{-i\omega\sqrt{2}\tau}\frac{\psi_{\omega l}^{*}(6M)}{6M}\,Y_{lm}^{*}\!\left(\frac{\pi}{2},\frac{\tau^{\prime}}{6\sqrt{3}M}\right)\,e^{i\omega\sqrt{2}\tau^{\prime}}\,.$ (16)
@@
[…]
×$\times\left|\sqrt{\frac{2l+1}{4\pi}\,\frac{(l-m)!}{(l+m)!}}\;P_{l}^{m}(0)\right|^{2}$
×$\times\exp\!\left[i\left(\frac{m}{6\sqrt{3}M}-\omega\sqrt{2}\right)\Delta\tau\right].$ (18)
@@ III.2 Derivation of the Excitation Rate $R(\Omega_{0})$
[…]
×$\times\int_{-\infty}^{\infty}d(\Delta\tau)\,e^{-i\Omega_{0}\Delta\tau}\,e^{i\left(\frac{m}{6\sqrt{3}M}-\omega\sqrt{2}\right)\Delta\tau},$ (19)
@@
×$\times\int_{0}^{\infty}d\omega\,\delta\!\left[\omega\sqrt{2}-\left(\frac{m}{6\sqrt{3}M}-\Omega_{0}\right)\right].$ (23)
@@ III.3 Numerical Results and Spectral Analysis
To elucidate the physical consequences of our derived transition rate, we now proceed with a numerical evaluation of Eq. (25). The analytical formula, while exact, contains an infinite sum over angular momentum modes and depends on the radial wave functions $\psi_{\omega_{m}l}(6M)$, which are not known in a simple closed form. To make progress, we employ the WKB approximation Chambers (1979) for the radial function, which is valid in the high-frequency regime $\omega^{2}\gg V_{l}(r)$ and provides the leading-order behavior as
@@ IV.1 Physical Interpretation of the Spectrum
The sharp, divergent nature of the peaks is a characteristic feature of first-order perturbation theory at resonance. In a more complete physical model, effects such as the natural lifetime of the excited state would regularize these divergences, broadening them into finite Lorentzian profiles. However, within our model, these sharp resonances powerfully signify the discrete, quantum nature of the atom’s interaction with the structured vacuum along its stable, periodic orbit. This unique spectral signature is a direct fingerprint of the kinematic and gravitational environment at the ISCO.
@@ IV.2 Comparison with Other Trajectories
The discrete, resonant spectrum predicted for an atom in a stable circular orbit is a unique signature of its specific trajectory. To fully appreciate its novelty, it is instructive to contrast this result with the well-established quantum phenomena experienced by atoms on other paths near a black hole. Our analysis, particularly the results visualized in Figure 2, demonstrates that the ISCO represents a special boundary case for stable orbital motion, producing a spectrum that is distinct from all other scenarios.
@@ IV.3 Potential implications and future work
From an astrophysical perspective, our results offer a new theoretical lens through which to view the dynamics of matter in the immediate vicinity of black holes. The innermost regions of accretion disks, culminating at the ISCO, are populated by ions and atoms subject to the extreme conditions we have modeled. The resonant excitation mechanism we have uncovered could, in principle, influence the emission spectra from these regions. The discrete frequencies at which the atom is preferentially excited might lead to corresponding bright or dark lines in the observed electromagnetic spectrum from the accretion disk. While in a realistic, dense plasma environment these sharp resonances would be significantly broadened, they could nonetheless contribute to the overall spectral features, providing a new quantum-level diagnostic for the dynamics near the ISCO.
@@
This study naturally points toward several compelling directions for future research: (1) the most immediate and critical next step is to extend this analysis to a rotating Kerr black hole. The vast majority of astrophysical black holes possess significant angular momentum, which dramatically alters the spacetime geometry, including the location and properties of the ISCO. The introduction of frame-dragging would undoubtedly lead to new spectral features, such as a splitting of the resonant peaks, providing a potential way to probe the black hole’s spin through quantum vacuum effects; (2) our use of a massless scalar field and a monopole detector serves as an essential first approximation. Future work should consider more realistic interactions, such as coupling the atom to the electromagnetic field. This would not only make the model more physically accurate but would also allow for the calculation of photon emission rates and polarization signatures, which are directly observable quantities; (3) a fascinating question is whether the excitation process generates entanglement. The related Unruh and Hawking effects are known to produce entangled particle pairs across horizons. Investigating the quantum state of the field after the atom’s excitation could reveal whether pairs of field quanta are emitted in an entangled state, with potential implications for relativistic quantum information; and (4) while our detector is a single quantum system, one could generalize this framework by considering an ensemble of orbiting atoms. By employing a quantum optics master equation approach, similar to the HBAR formalism, one could study the collective emission and explore the thermodynamics of the resulting radiation field, investigating whether a steady-state thermal distribution can be achieved and what its properties might be.
@@ V Conclusion
Acknowledgements.
R. P. and A. Ö would like to acknowledge networking support of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG), and the COST Action CA23115 - Relativistic Quantum Information (RQI) funded by COST (European Cooperation in Science and Technology). R. P. and A. Ö would also like to acknowledge the funding support of SCOAP3.
Data Availability Statement

### v2

@@ I Introduction
The synthesis of quantum mechanics and general relativity remains one of the most profound challenges in modern theoretical physics Rovelli_2004. In this pursuit, the study of quantum fields in curved spacetimes has emerged as an indispensable tool, offering a window into the fundamental interactions that govern our universe Birrell:1982ix. Black holes, as cosmic laboratories of extreme gravity, provide a unique arena to probe the intricate link between quantum phenomena and the fabric of spacetime itself Bardeen:1973gs. The behavior of quantum fields in the vicinity of these enigmatic objects not only deepens our understanding of foundational physics but also holds direct relevance for astrophysical observables, such as the celebrated Hawking radiation and the complex dynamics of accretion disks Page:1974he; Thorne:1974ve; Frank_2002.
@@
A particularly fascinating frontier in this domain is the interaction between localized quantum systems, such as atoms, and the quantum fields permeating the curved background of a black hole. When an atom is subjected to a strong gravitational field, its internal energy levels are perturbed not only by the spacetime curvature but also by the incessant fluctuations of the quantum vacuum Audretsch:1995qb. This interaction can lead to observable phenomena, providing a novel way to probe the structure of both the gravitational field and the quantum vacuum itself. The theoretical underpinnings for such investigations are well-established, rooted in the Unruh effect, which predicts that an accelerated observer in flat spacetime will perceive the vacuum as a thermal bath, and its gravitational analogue, Hawking radiation, where black holes emit thermal particles due to quantum effects near their event horizon Hawking:1975vcx; Unruh:1976db; Takagi_1986.
@@
Building on these foundational concepts, previous studies have extensively examined the response of Unruh-DeWitt detectors, which are simplified yet powerful models for atoms, on various trajectories, including those held at static positions or undergoing radial infall toward a black hole. More recently, a quantum optics approach has been successfully applied to investigate the ”horizon brightened acceleration radiation” (HBAR) emitted by atoms falling into a black hole, revealing a deep connection between acceleration radiation, thermodynamics, and near-horizon conformal symmetries Scully:2017utk; Azizi:2021qcu; Azizi:2021yto; Ordonez:2025sqp; Bukhari:2023yuy; Bukhari:2022wyx; Ovgun:2025isv; Pantig:2025okn; Das:2025rzz.. This framework has proven to be a versatile tool for analyzing such complex quantum interactions.
@@
While these studies have provided invaluable insights, the specific case of a quantum detector following a stable circular geodesic has received comparatively less attention Garay:1999sk; Louko:2006zv; Casals:2009zh; Hodgkinson:2014iua; Ng:2014kha; Biermann:2020bjh. This paper seeks to fill this gap by focusing on a region of profound astrophysical and theoretical importance: the ISCO. Located at a radius of ($r=6M$) for a Schwarzschild black hole, the ISCO represents the final frontier of stable orbital motion, a critical boundary where matter in accretion disks accumulates before its final plunge. Understanding quantum interactions at this precipice is multifaceted Ottewill:2000qh. It promises to shed light on how quantum systems behave under extreme gravitational acceleration, potentially revealing novel quantum-gravitational effects. Furthermore, an atom orbiting at the ISCO can act as a sensitive probe of the quantum vacuum structure, offering a way to study effects akin to Hawking radiation without the need for direct observation of the horizon itself Barbado:2011dx.
@@ II.1 Schwarzschild Spacetime and Geodesics at the ISCO
The gravitational field of a non-rotating, uncharged, spherically symmetric black hole of mass $M$ is described by the Schwarzschild metric. In geometric units, where the gravitational constant $G$ and the speed of light $c$ are set to unity, the line element in Schwarzschild coordinates $(t,r,\theta,\varphi)$ is given by Carroll_2019
@@ II.2 The Unruh-DeWitt Detector and the Quantum Field
where $\sigma^{+}=|e\rangle\langle g|$ and $\sigma^{-}=|g\rangle\langle e|$ are the standard atomic raising and lowering operators, respectively. The interaction between the detector and a massless scalar field $\phi$ is described by the interaction Hamiltonian, written in the interaction picture as Unruh:1976db; DeWitt_1979
@@
$\phi(x)=\sum_{l=0}^{\infty}\sum_{m=-l}^{l}\int_{0}^{\infty}d\omega\left[a_{\omega lm}u_{\omega lm}(x)+a^{\dagger}_{\omega lm}u^{*}_{\omega lm}(x)\right],$ (11)
@@
The choice of vacuum state, $|0\rangle$, is critical as it defines the initial state of the field with which the atom interacts. For an observer located at asymptotic infinity, the natural vacuum state is the Boulware vacuum, denoted $|0_{\text{B}}\rangle$. This state is defined by the condition that it is annihilated by the operators $a_{\omega lm}$ corresponding to mode functions $u_{\omega lm}$ that are purely positive frequency with respect to the global timelike Killing vector $tial_{t}$. Physically, the Boulware vacuum corresponds to a state with no particles and no radiation as seen by a static observer far from the black hole. This is the ideal choice for our analysis, as it allows us to isolate the excitation of the atom that is due solely to its orbital motion, ensuring that our results are not conflated with the thermal bath of particles associated with the Hawking effect, which is characteristic of other vacuum states like the Hartle-Hawking vacuum. The orbiting atom is thus treated as a local accelerated probe of this global, non-thermal vacuum state.
@@ III.1 The Wightman Function along the ISCO Trajectory
The transition rate $R(\Omega_{0})$ for a detector with energy gap $\Omega_{0}$ is given by Fermi’s Golden Rule Brown_1992. For a stationary trajectory, such as a circular orbit, it is expressed as the Fourier transform of the Wightman function with respect to the proper time difference $\Delta\tau=\tau-\tau^{\prime}$:
@@
$G^{+}(x,x^{\prime})=\sum_{l,m}\int_{0}^{\infty}d\omega\,u_{\omega lm}(x)u_{\omega lm}^{*}(x^{\prime}),$ (14)
@@
Here, $\psi_{\omega l}(r)$ are the radial functions that solve the Regge-Wheeler equation, and $Y_{lm}(\theta,\varphi)$ are the standard spherical harmonics. Our task is to evaluate the Wightman function specifically along the atom’s ISCO worldline, $x^{\mu}(\tau)=(\sqrt{2}\tau,6M,\pi/2,\tau/6\sqrt{3}M)$. Substituting these coordinates into the mode functions, we obtain the Wightman function as experienced by the orbiting atom as
$G^{+}(\tau,\tau^{\prime})$ $=\sum_{l,m}\int_{0}^{\infty}d\omega\,\frac{\psi_{\omega l}(6M)}{6M}\,Y_{lm}\!\left(\frac{\pi}{2},\frac{\tau}{6\sqrt{3}M}\right)$
$\times e^{-i\omega\sqrt{2}\tau}\frac{\psi_{\omega l}^{*}(6M)}{6M}\,Y_{lm}^{*}\!\left(\frac{\pi}{2},\frac{\tau^{\prime}}{6\sqrt{3}M}\right)\,e^{i\omega\sqrt{2}\tau^{\prime}}\,.$ (16)
@@
$G^{+}(\Delta\tau)$ $=\sum_{l,m}\int_{0}^{\infty}d\omega\,\frac{|\psi_{\omega l}(6M)|^{2}}{(6M)^{2}}$
$\times\left|\sqrt{\frac{2l+1}{4\pi}\,\frac{(l-m)!}{(l+m)!}}\;P_{l}^{m}(0)\right|^{2}$
$\times\exp\!\left[i\left(\frac{m}{6\sqrt{3}M}-\omega\sqrt{2}\right)\Delta\tau\right].$ (18)
@@ III.2 Derivation of the Excitation Rate $R(\Omega_{0})$
$R(\Omega_{0})$ $=g^{2}\sum_{l,m}\frac{|C_{lm}|^{2}}{(6M)^{2}}\int_{0}^{\infty}d\omega\,|\psi_{\omega l}(6M)|^{2}$
$\times\int_{-\infty}^{\infty}d(\Delta\tau)\,e^{-i\Omega_{0}\Delta\tau}\,e^{i\left(\frac{m}{6\sqrt{3}M}-\omega\sqrt{2}\right)\Delta\tau},$ (19)
@@
$\times\int_{0}^{\infty}d\omega\,\delta\!\left[\omega\sqrt{2}-\left(\frac{m}{6\sqrt{3}M}-\Omega_{0}\right)\right].$ (23)
@@ III.3 Numerical Results and Spectral Analysis
To elucidate the physical consequences of our derived transition rate, we now proceed with a numerical evaluation of Eq. (25). The analytical formula, while exact, contains an infinite sum over angular momentum modes and depends on the radial wave functions $\psi_{\omega_{m}l}(6M)$, which are not known in a simple closed form. To make progress, we employ the WKB approximation Chambers_1979; Schutz:1985km for the radial function, which is valid in the high-frequency regime $\omega^{2}\gg V_{l}(r)$ and provides the leading-order behavior as
@@ IV.1 Physical Interpretation of the Spectrum
The sharp, divergent nature of the peaks is a characteristic feature of first-order perturbation theory at resonance. In a more complete physical model, effects such as the natural lifetime of the excited state would regularize these divergences, broadening them into finite Lorentzian profiles. However, within our model, these sharp resonances powerfully signify the discrete, quantum nature of the atom’s interaction with the structured vacuum along its stable, periodic orbit. This unique spectral signature is a direct fingerprint of the kinematic and gravitational environment at the ISCO. It is important to note that this discrete structure arises from the worldline’s periodicity and is not an artifact of the Boulware vacuum. While a different choice of vacuum (e.g., a thermal state) would alter the background and the peak amplitudes, the resonant frequency comb, which is a purely kinematic effect, would persist.
This resonant phenomenon is not unique to gravitational contexts but has direct analogues in other areas of physics, such as cavity optomechanics and trapped-ion physics Aspelmeyer:2013fha. In cavity optomechanics, the periodic motion of a mechanical oscillator coupled to an optical cavity mode leads to the creation of motional sidebands in the optical spectrum, a process in which mechanical phonons are converted into optical photons. The underlying physics is the same: a periodic modulation of the interaction between a quantum system and a field opens up discrete frequency channels for energy exchange. In our case, the atom’s orbital motion plays the role of the mechanical oscillator, and its excitation corresponds to the creation of a quantum from the vacuum field, with the allowed energies dictated by the harmonics of the orbital frequency. This analogy reinforces the interpretation that the predicted frequency comb is a robust physical consequence of periodic motion.
@@ IV.2 Comparison with Other Trajectories
The discrete, resonant spectrum predicted for an atom in a stable circular orbit is a unique signature of its specific trajectory. Previous studies have indeed established that detectors on general circular geodesics experience a non-thermal vacuum response that is distinct from the thermal spectrum seen by a static, accelerated observer Hodgkinson:2014iua; Ng:2014kha; Biermann:2020bjh. Our work builds on this foundation by providing a novel characterization of this phenomenon in several key respects. Firstly, we are the first to investigate the quantum signature of the ISCO, a fundamental physical boundary of profound astrophysical interest. Secondly, we explicitly identify the spectrum’s structure as a ”frequency comb,” where the teeth are harmonics of the proper orbital frequency. Most importantly, our analysis of the approach to the ISCO, as shown in Fig. 2, reveals a dramatic amplification of the excitation rate and a widening of the peak spacing, rendering the ISCO’s spectral signature qualitatively distinct from that of any other stable orbit. To fully appreciate its novelty, it is instructive to contrast this result with the quantum phenomena experienced by atoms on other paths near a black hole. Our analysis, particularly the results visualized in Figure 2, demonstrates that the ISCO represents a special boundary case for stable orbital motion, producing a spectrum that is distinct from all other scenarios.
@@ IV.3 Potential implications and future work
From an astrophysical perspective, our results offer a new theoretical lens through which to view the dynamics of matter in the immediate vicinity of black holes. The innermost regions of accretion disks, culminating at the ISCO, are populated by ions and atoms subject to the extreme conditions we have modeled. The resonant excitation mechanism we have uncovered could, in principle, influence the emission spectra from these regions. The discrete frequencies at which the atom is preferentially excited might lead to corresponding bright or dark lines in the observed electromagnetic spectrum from the accretion disk. While in a realistic, dense plasma environment, these sharp resonances would be significantly broadened, they could nonetheless contribute to the overall spectral features, providing a new quantum-level diagnostic for the dynamics near the ISCO. Observationally, the signature of this effect would likely not be sharp lines but rather a series of quasi-periodic oscillations (QPOs) in the disk’s high-frequency emission spectrum. The key to filtering this signal from the bright thermal background would be to search for the characteristic harmonic structure predicted by our model. A time-series analysis revealing a ”comb” of QPO frequencies that are integer multiples of a fundamental frequency tied to the ISCO orbital period would provide a distinct fingerprint of this quantum resonant mechanism, distinguishing it from other phenomena.
@@
This study naturally points toward several compelling directions for future research: (1) the most immediate and critical next step is to extend this analysis to a rotating Kerr black hole. The vast majority of astrophysical black holes possess significant angular momentum, which dramatically alters the spacetime geometry, including the location and properties of the ISCO. The introduction of frame-dragging would undoubtedly lead to new spectral features, such as a splitting of the resonant peaks, providing a potential way to probe the black hole’s spin through quantum vacuum effects Starobinsky:1973a; (2) our use of a massless scalar field and a monopole detector serves as an essential first approximation, successfully isolating the primary kinematic effects of the orbit. A crucial next step is to consider more realistic interactions by coupling the atom to the electromagnetic field via its electric dipole moment. This introduces significant new physics, including polarization dependence and atomic selection rules. We have outlined the mathematical framework for this extension in Appendix A. The full calculation, which requires the evaluation of the electromagnetic field correlator in curved spacetime, remains a compelling direction for future work. (3) a fascinating question is whether the excitation process generates entanglement. The related Unruh and Hawking effects are known to produce entangled particle pairs across horizons. Investigating the quantum state of the field after the atom’s excitation could reveal whether pairs of field quanta are emitted in an entangled state, with potential implications for relativistic quantum information; and (4) while our detector is a single quantum system, one could generalize this framework by considering an ensemble of orbiting atoms. By employing a quantum optics master equation approach, similar to the HBAR formalism, one could study the collective emission and explore the thermodynamics of the resulting radiation field, investigating whether a steady-state thermal distribution can be achieved and what its properties might be.
@@ V Conclusion
Acknowledgements R. P. and A. Ö would like to acknowledge networking support of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers), the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES), the COST Action CA21136 - Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse), the COST Action CA23130 - Bridging high and low energies in search of quantum gravity (BridgeQG), and the COST Action CA23115 - Relativistic Quantum Information (RQI) funded by COST (European Cooperation in Science and Technology). R. P. and A. Ö would also like to acknowledge the funding support of SCOAP3.
VI Data Availability Statement